# What’s Wrong with Your Code Generated by Large Language Models? An Extensive Study

### v1

@@ 5.2. Effectiveness in RWPB and Bugs Analysis
We analyzed the bug distribution in the real-world project benchmark, RWPB and compared the distribution with HumanEval+, MBPP+, and APPS+. Similar to the existing benchmarks, the proportion of syntax bugs in RWPB is the lowest, indicating that LLMs can effectively avoid syntax bugs in real-world scenarios as well. Furthermore, Figure 2 demonstrates that the comment number between correct and incorrect code in RWPB exhibits the same trend as the existing benchmarks: incorrect code contains more comments than correct code. What’s more, the distribution of misunderstanding and logic error is similar. Figure 4 shows that missing checks for corner cases are the most prevalent sub-categories (53.2%), followed by the residual logic misunderstanding (30.4%). However, unlike existing benchmarks, the proportion of runtime bugs is higher for all LLMs, particularly for incorrect arguments and incorrect boundary condition checks.
@@
Finding 12: For RWPB, LLMs align with existing benchmarks in most parts. However, the proportion of runtime bugs, especially incorrect arguments and incorrect boundary condition checks, is significantly higher compared to existing benchmarks.
We further analyzed the incorrect code with incorrect arguments. We found that, different from existing benchmarks, real-world projects contain a richer variety of argument types, such as multi-dimensional tensors. As a result, generated code often fails due to dimension mismatches in these arguments. This suggests that LLMs need to perform more effective inference on parameters to avoid incorrect arguments in real-world projects.
Finding 13: RWPB involves diverse argument types, leading high proportion of incorrect arguments.
In real-world projects, handling high-dimensional variables such as Tensor arrays often requires extensive boundary checks across multiple dimensions. For instance, consider the following code that generates and fills a mask to fit a given context window. However, it neglects boundary checks for the token dimension. When the token is a one-dimensional array, it can lead to out-of-bounds access issues.
import numpy as np
def pad_tokens(tokens, context_window, pad_token):
max_seq_len = np.max(tokens.shape)
# tuple index out of range
mask = np.arange(max_seq_len)[None, :] < tokens.shape[1]
Finding 14: For RWPB that involves high-dimensional variables, it is crucial for the LLM to perform boundary checks on dimension information to enhance code robustness.
6. Mitigating Bugs in Generated Code
In this section, we discuss how LLMs identify and fix bugs in their incorrect code in the inference time.
Our proposed method is based on self-critique and is cost-effective. Self-critique is the process by which LLMs critique and review the responses they generate to enhance their understanding and reasoning ability in problem-solving (Gero et al., 2023; Gou et al., [n. d.]; Shinn et al., 2023; Tan et al., 2023; Saunders et al., 2022). Wang et al. (Tan et al., 2023) introduced self-criticism during the fine-tuning phase to reduce the harmful responses generated by LLMs. Saunders et al. (Saunders et al., 2022) fine-tuned LLMs to generate natural language critiques to help humans find flaws in a topic-based summarization task. Inspired by these methods, we introduce self-critique in the inference time, making LLMs critique the code, and then fix these bugs. Specifically, we re-design the prompt template to contain the original problem description, incorrect code, and our constructed bug category, making LLMs analyze the root causes of bugs and identify locations of bugs in the code. Based on the root causes and locations of bugs, we then make them re-correct the code they generated. To help LLMs better understand the problem and original error code, we also incorporate the feedback of incorrect code from the compiler (i.e., error messages from failed executions) within the prompt template. The whole process is fully automated.
@@ 6. Mitigating Bugs in Generated Code
Figure 6. The prompt template of LLMs to critique and re-correct the incorrect code.
@@
The simplified prompt for critiquing and fixing is shown in Figure 6. \$DOC_STRING and \$INCORRECT_CODE denote the original programming problem and the incorrect code generated by LLM. \$BUG_CATEGORY contains our constructed bug category and the natural interpretation of bug types. \$COMPILER_FEEDBACK is the outcome of the compiler execution. If the result belongs to type A (i.e., Syntax Bug) and type B (i.e., Runtime Bug), we provide the error log; if the compiler does not return an error but the code does not pass the unit tests (i.e., Functional Bug), we provide the message “The functionality of code is incorrect.”; and if the execution of the code exceeds the time limit, we offer the message “The execution of the code has exceeded the time limit.” \$CAUSE_AND_CODE denotes the root causes and locations of bugs, and the re-corrected code generated by LLM. The final version of the prompts is available in our artifact repository (ar2, 2023). The LLM used for critiquing code and fixing bugs is the same model utilized for generating code based on human requirements. Consequently, our method eliminates the need for additional post-training costs.
@@
Our proposed method is interactive and iterative. We extract the re-corrected code from \$CAUSE_AND_CODE through scripts and then verify the correctness of the code by interacting with the compiler. If the re-corrected code still does not pass unit tests, we update the prompt template with the latest incorrect code and the compiler outcome to conduct the next re-correct iteration.
Figure 7. Results of fixing bugs by using our self-critic-based approach.
@@
Effectiveness of our proposed method in mitigating bugs. To evaluate the effectiveness of our fixing bugs method, we conducted extensive experiments across the prior seven LLMs. For each LLM, we randomly selected 120 incorrect codes generated by the respective models, then utilized our methods of self-critique and self-revision on the code to fix bugs. We conducted two iterations of our method to fix bugs on each incorrect code. The performance on GPT-4 is shown in Figure 7 and results for the other six LLMs are available in our artifact repository (ar2, 2023).) The results suggest that our proposed method can effectively identify and fix bugs in incorrect code, thereby enhancing the success rate of code generation using LLMs. Specifically, in the initial iteration, GPT-4 fixes 24.1% (29 in 120) of incorrect self-generated code through the proposed method. A further 6.6% (6 in 91) of the codes are corrected in the second iteration. The syntax bugs are completely fixed in the first iteration. For runtime bugs, 23% of error codes are corrected in the first iteration and 16% in the second iteration. For functional bugs, 23% are fixed in the first iteration and 3% in the second iteration.
@@
Figure 7 also presents the evolution of bugs for GPT-4. For runtime bugs, 46% and 32% are transformed to functional bugs at the end of the first and second iterations, respectively. Conversely, the functional bugs are harder to transform to the two other types of bugs. 22% and 4% functional bugs are fixed and only 8% and 3% bugs transformed to runtime bugs after two iterations, respectively. Experimental results show that our approach can effectively fix all types of code to improve the performance of LLMs in code generation, while facilitating the transition of syntax and runtime bugs towards functional bugs.
Additionally, although experiments demonstrate that our method improves LLM code generation through self-critique, further enhancements can be pursued. From an artificial intelligence standpoint, introducing comments in code training corpora and employing few-shot prompt learning during the inference phase could enhance code quality. From a software engineering perspective, incorporating compiler feedback with post-training could help LLMs better understand code-related tasks.
7. Discussion
Benchmark Analysis. HumanEval+ and MBPP+ are widely used benchmarks in code generation with high-quality problem descriptions, diverse unit tests, and formatted input/output. The APPS+, sourced from various algorithm online-judge websites, presents high-difficulty problems. However, algorithmic problems result in lengthy and often redundant descriptions (e.g., extensive background to introduce the problem), complicating comprehensive understanding by LLMs. Furthermore, the algorithmic problems originally included images as the problem explanation, but these images are missing in the APPS+, further increasing the difficulty for models to understand the tasks. To ensure experiment quality, we manually excluded problems with missing images.
Annotation Process. During the process of analyzing and labeling bug types in incorrect code, experts are allowed to utilize GPT-4 to aid in annotation tasks. Specifically, for functional bugs that lack detailed compiler error information, we employ GPT-4 to generate three potential root causes with explanations, facilitating the analysis and classification of bug types. Although the final annotations are manually conducted and subject to double-checking, we are interested in exploring the differences in effectiveness and efficiency among three annotation methods: fully automated, fully manual, and AI-assisted, in identifying bug types in generated code.
We conduct an experiment to evaluate different annotation methods for bug types in incorrect code. We randomly select 100 incorrect codes and engage five experts who have not labelled these codes, to manually annotate the bug types separately without AI assistance. Another group of five experts annotates the same codes, but with the assistance of GPT-4, which provides potential root causes for the bugs. Additionally, we incorporate a bug taxonomy in the prompt to guide GPT-4 in analyzing and classifying the bugs according to the taxonomy. For each annotation method, we compare the average result with the double-checking results. When annotating tertiary categories, the accuracy of fully manual, AI-assisted and fully automated are 91.7%, 84.4% and 5.2%, respectively. Results suggest that AI-assisted annotation can improve the accuracy of annotation compared to fully manual annotation, which is a conclusion similar to the findings of Saunders et al. (Saunders et al., 2022) on the topic-based summarization task. Moreover, GPT-4 is not yet able to replace the manual annotation to achieve fully automated annotation in our task. Furthermore, we evaluate the accuracy of fully automated annotation on secondary categories, which stands at 35.7%. This indicates that GPT-4 can provide coarse-grained bug categories of bugs to aid expert annotation. On the other hand, AI-assisted can enhance the efficiency of the annotation process. The experts provide feedback that annotating with the help of bug causes given by the GPT-4 can expedite their labelling tasks.
During the annotation process, we also observe that a small proportion of annotators heavily rely on the root causes provided by GPT-4, neglecting manual analysis of the bugs in the incorrect code. We refer to this behavior as “lazy annotation”. To mitigate this issue, we implement a strategy where 10% of the annotations are selected for double-checking manually, and we emphasize the importance of manual error code analysis to the annotators. This phenomenon of “lazy annotation” appears to be common across AI-assisted labeling scenarios and presents a challenge to maintaining high-quality annotations. In the future, we plan to further investigate methods to prevent “lazy annotation” in various annotation tasks to enhance both the efficiency and effectiveness of the labeling process.
8. Threats to Validity
@@
In this section, we discuss some significant threats to the validity of our results.

### v2

@@ 5.2 Effectiveness in RWPB and Bugs Analysis
Bug Analysis. We analyzed the bug distribution in the real-world project benchmark, RWPB, and compared the distribution with HumanEval+, MBPP+, and APPS+. Similar to the existing benchmarks, the proportion of syntax bugs in RWPB is the lowest, indicating that LLMs can effectively avoid syntax bugs in real-world scenarios as well. Furthermore, Figure 2 demonstrates that the number of comments between correct and incorrect code in RWPB exhibits the same trend as the existing benchmarks: incorrect code contains more comments than correct code. What is more, the distribution of misunderstanding and logic errors is similar. Figure 4 shows that missing checks for corner cases are the most prevalent sub-categories (53.2%), followed by the residual logic misunderstanding (30.4%). However, unlike existing benchmarks, the proportion of runtime bugs is higher for all LLMs, particularly for incorrect arguments and incorrect boundary condition checks.
@@
Finding 9: For RWPB, LLMs align with existing benchmarks in most parts. However, the proportion of runtime bugs, especially incorrect arguments and incorrect boundary condition checks, is significantly higher compared to existing benchmarks. Implication: Encouraging the model to perform self-critique, such as checking the input format aligned with the requirement, can improve its performance.
6 Mitigating Bugs in Generated Code
In the preceding section, we analyzed and categorized the types of bugs in the LLM-generated code, yielding several interesting findings. Building on these insights, we propose a straightforward and effective method that enables LLMs to self-critique their generated code and iteratively fix bugs.
6.1 Method
Recent studies [68, 69] have demonstrated that providing LLMs with more detailed information about specific tasks can enhance their performance on those tasks, such as in math reasoning [70]. Additionally, LLMs can leverage the world knowledge acquired during the pre-training phase to critique text, aiding humans in identifying errors in natural language. For instance, Saunders et al. [71] fine-tuned LLMs to generate natural language critiques to help humans find flaws in a topic-based summarization task. Wang et al. [46] also introduced critique data in the fine-tuning phase to reduce the harmful responses generated by LLMs. Inspired by these findings, we introduce a self-critique method to enable LLMs to analyze and critique their own generated code, to identify hidden bugs, and subsequently correct them.
@@ 6.1 Method
Figure 6: The prompt template of LLMs to critique and re-correct the incorrect code.
In practice, we found that current LLMs that have undergone supervised fine-tuning (SFT) can already analyze and critique effectively. Therefore, we utilize in-context learning to mitigate bugs directly, rather than creating additional data for them to learn how to critique. We design a prompt template that enables LLMs to critique and reflect on their generated code according to the programming problem and the incorrect code, identifying the locations of bugs within the code. Based on the root causes and locations of these bugs, we then prompt the LLMs to correct the generated code. Additionally, we incorporate our constructed bug taxonomy and feedback from the compiler (i.e., error messages from failed executions) within the prompt template, providing both holistic and individual information. This helps LLMs better understand the problem and the original error code, to further improve the accuracy and performance of the critique process. The entire process is fully automated.
@@
The simplified prompt for our method is shown in Figure 6. \$DOC_STRING denotes the original programming problem and \$INCORRECT_CODE denotes the incorrect code generated by LLMs. \$BUG_CATEGORY contains our constructed bug category and the natural interpretation of bug types. \$COMPILER_FEEDBACK is the outcome of the compiler execution. If the result belongs to type A (i.e., Syntax Bug) and type B (i.e., Runtime Bug), we provide the compiler bug report; if the compiler run successfully but the code does not pass the unit tests (i.e., Functional Bug), we provide the message “The functionality of code is incorrect”; and if the execution of the code exceeds the time limit, we offer the message “The execution of the code has exceeded the time limit”. \$CAUSE_AND_CODE denotes the root causes and locations of bugs, and the re-corrected code generated by LLM. The final version of the prompts is available in our artifact repository [54]. The LLM used for critiquing code and fixing bugs is the same model utilized for generating code based on human requirements. Consequently, our method eliminates the need for additional post-training costs.
@@
We extract the re-corrected code from \$CAUSE_AND_CODE through scripts and then verify the correctness of the code by interacting with the compiler. If the re-corrected code still does not pass unit tests, we update the prompt template with the latest incorrect code and the compiler outcome to conduct the next re-correct iteration. In summary, our proposed self-critique-based method is cost-effective, interactive, and iterative.
6.2 Evaluation
6.2.1 Effectiveness
@@
To evaluate the effectiveness of our method, we conducted extensive experiments across the prior nine LLMs. For each LLM, we randomly selected 120 incorrect codes generated by the respective models, then utilized our methods of self-critique and self-revision on the code to fix bugs. We conducted n iterations (i.e., we set n = 2 in this experiment) of our method to fix bugs in each incorrect code. The performance on GPT-4 is shown in Figure 7. The results suggest that our proposed method can effectively identify and fix bugs in incorrect code, thereby enhancing the success rate of code generation using LLMs. Specifically, in the first iteration, GPT-4 fixes 24.1% (29 in 120)) of incorrect self-generated code through the proposed method. A further 6.6% (6 in 91) of the codes are corrected in the second iteration. Therefore, our method achieves a repair success rate of 29.2% (i.e., (29+6)/120 = 29.2%) after two iterations.
The syntax bugs are completely fixed in the first iteration. For runtime bugs, 23% of error codes are corrected in the first iteration and 16% in the second iteration. For functional bugs, 23% are fixed in the first iteration and 3% in the second iteration. The additional experimental results for the other eight LLMs are available in our artifact repository [54].
6.2.2 Comparison Results
To further demonstrate the effectiveness of our proposed bug repair framework, we also compare it with two popular code repair methods, i.e., Self-debug [72] and Intervenor [73]. Self-debug employs prompt engineering to guide the LLM in repairing bugs based on compiler feedback. Intervenor first prompts the teacher model to generate bug repair suggestions (chain-of-thought) based on compiler feedback, and then the student model applies these suggestions to modify the code. All three approaches were evaluated under identical experimental settings, using the same test data and GPT-4, as shown in Table 6.
Table 6: Comparison of methods for repairing code bugs. BT. Denotes Bug Taxonomy.
Method Repair Success Rate
Iteration 1 Iteration 2
Self-debug [72] 23.3% 27.5%
Intervenor [73] 22.5% 28.3%
Ours 24.1% 29.2%
   w/o BT. 21.7% 27.5%
Ours + Intervenor 25.8% 29.2%
Experimental results indicate that our framework outperforms both self-debug and Intervenor. The self-debug approach achieves a lower repair success rate because it relies solely on compiler feedback for modifications. Although Intervenor introduces a teacher model to generate repair suggestions and a student model to apply code changes, it essentially remains dependent on compiler feedback for both explanation and modification. In comparison, our framework provides additional information beyond compiler results, including the potential bug type and a detailed bug explanation, which helps the LLM more accurately identify and fix bugs.
To further validate the effectiveness of our findings, we incorporated bug classification information into Intervenor’s teacher model to assess whether this additional information could improve the accuracy of repair suggestions (i.e., introducing our approach’s insights into the Intervenor framework). The results, presented as “Ours + Intervenor,” demonstrate that our findings provide supplementary repair information, enabling Intervenor to fix more bugs. In summary, both the findings and our framework are effective in enhancing the LLM’s understanding of bugs and facilitating bug repair.
6.2.3 Ablation Study
To further validate the effectiveness of bug taxonomy in our bug repair framework, we conducted ablation experiments by comparing the bug repair success rate with and without providing the bug taxonomy in the prompt. To ensure a fair comparison, we adopted the same experimental settings. The results are presented in Table 6. From experimental results, we draw several conclusions. First, the absence of bug taxonomy in the prompt led to a significant decrease in the code repair success rate. Without the bug taxonomy, the repair methods degenerate to relying solely on compiler feedback for bug localization. In particular, for type 3 functional bugs (constitute a significantly larger proportion of bugs, as shown in Tables 3, 4 and 5), compiler feedback alone provides limited information, typically indicating only whether the code is correct or incorrect. Therefore, supplementing the prompt with a checklist of possible Python bug types helps LLMs systematically consider potential errors.
Notably, when the bug taxonomy is omitted, all three methods, including self-debug [72] and Intervenor [73], rely only on compiler feedback for bug repair. However, differences in prompt design still result in performance variations among these methods. Finally, across two iterations, the bug repair pipelines that included the bug taxonomy consistently fixed more bugs than those without it.
Figure 7: Results of fixing bugs by using our self-critical-based approach.
6.2.4 Bug Evolution
@@
Figure 7 also presents the bug evolution for GPT-4. For runtime bugs, 46% and 32% are transformed to functional bugs at the end of the first and second iterations, respectively. Conversely, the functional bugs are harder to transform into the two other types of bugs. 22% and 4% functional bugs are fixed, and only 8% and 3% bugs are transformed into runtime bugs after two iterations, respectively. Experimental results show that our approach can effectively fix all types of code to improve the performance of LLMs in code generation, while facilitating the transition of syntax and runtime bugs towards functional bugs.
In summary, in this section, we propose a bug repair framework that enables LLMs to fix their generated bugs using a self-critique mechanism. Unlike two iterative bug repair methods we compared, our framework introduces the bug taxonomy and bug type explanations distilled from our analysis and findings. Notably, this framework primarily serves as a straightforward application scenario for the main findings and contributions of our study. Its purpose is to demonstrate how one aspect of our research, i.e., our bug taxonomy and bug type explanations, can be applied to practical bug repair, thereby highlighting the value of our work.
Building on our study, further enhancements can be explored. For example: (1) The bug taxonomy we constructed can serve as a valuable foundation for reward function design in reinforcement learning, with the potential to improve the effectiveness of reinforcement learning from compiler feedback (RLCF) approaches. (2) Moreover, we found that incorporating information about code comments and function parameters significantly improves LLMs’ understanding of functions in real-world projects. Providing such information during training or at inference time may further enhance the code generation capabilities of LLMs. (3) In Section 5 and Section 7, we detail how we automated the iterative synthesis of unit tests and function docstrings during dataset construction, and how dataset quality was ensured through manual double-checking by annotators. This process provides important references for future researchers aiming to automate the construction of other types of code datasets. We hope our findings provide the community with deeper insights into LLM code generation and contribute to the development of more robust models and more efficient code repair techniques.
7 Discussion
@@
In this section, we discuss some findings during the experiments, future work, and some significant threats to the validity of our results.
7.1 Benchmark Analysis
HumanEval+ and MBPP+ are widely used benchmarks in code generation with high-quality problem descriptions, diverse unit tests, and formatted input/output. APPS+, sourced from various algorithm online-judge websites, presents high-difficulty problems. However, algorithmic problems result in lengthy and often redundant descriptions (e.g., extensive background to introduce the problem), complicating comprehensive understanding by LLMs. Furthermore, the algorithmic problems originally included images as the problem explanation, but these images are missing in the APPS+, further increasing the difficulty for models to understand the tasks. To ensure experiment quality, we manually excluded problems with missing images.
7.2 Annotation Process
During the process of analyzing and labeling bug types in incorrect code, experts are allowed to utilize GPT-4 to aid in annotation tasks. Specifically, for functional bugs that lack detailed compiler error information, we employ GPT-4 to generate three potential root causes with explanations, facilitating the analysis and classification of bug types. Although the final annotations are manually conducted and subject to double-checking, we are interested in exploring the differences in effectiveness and efficiency among three annotation methods: fully automated by GPT-4, fully manual, and AI-assisted, in identifying bug types.
We randomly select 100 incorrect code samples and engage five experts who have not previously labeled these codes to separately and manually annotate the bug types without AI assistance. Another group of five experts uses GPT-4 for assistance to separately annotate bugs in the same manner as before. For fully automated annotation, we provide GPT-4 with the compiler results (if have) and the bug taxonomy, and task it with analyzing the incorrect code and classifying the bug. The temperature and top-k are 0.1 and one, respectively. We take the double-checked results as ground truths and report the average accuracy of these three annotation methods, as shown in Table 7. The results show that, on secondary categories, the accuracy of fully automated annotation stands at 37%. This indicates that GPT-4 can provide coarse-grained bug categories of bugs to aid expert annotation. For tertiary categories, the accuracy of AI-assisted, fully manual, and fully automated is 91.6%, 84.4%, and 5%, respectively. It indicates that AI-assisted annotation enhances accuracy compared to fully manual methods, which is a conclusion similar to the findings of Saunders et al. [71] on the topic-based summarization task. Moreover, fully automated annotation is not yet able to replace manual annotation to achieve fully automated annotation in our task. On the other hand, AI-assisted can enhance the efficiency of the annotation process. The experts provide feedback that annotating with the help of bug causes given by the GPT-4 can expedite their labeling tasks.
Table 7: The average accuracy of three annotation methods (i.e., AI-assisted, fully manual, and fully automated) in identifying bug types.
Bug Category Level AI-Assisted Fully Manual Fully Automated
Secondary Category 98.2% 99.0% 37%
Tertiary Category 91.6% 84.4% 5%
7.3 Automating RWPB Construction Process
In Section 5.1, we propose a pipeline to manually construct a real-world benchmark, while trying our best to prevent data leakage. However, an ideal situation would be to automate the construction of high-quality datasets. The automated process can easily synthesize large amounts of data that can be used not only for evaluation, but also to expand the size of our training data, increasing the ease of use of the method. In this section, we discuss the strategy for automating our manual construction process.
The initial RWPB build process can be summarized into five parts: (1) start with real-world code repositories, filtering for those that are recently released and popular; (2) use automated scripts to filter and clean the functions extracted from these repositories; (3) remove duplicate functions using both open-source and closed-source large-scale code corpora; (4) manually synthesize docstrings, including programming problem descriptions, arguments, and returns; (5) generate unit test examples and ensure they cover the entire code comprehensively. In fact, the first, second, and third steps, as well as the calculation of unit test coverage in the fifth step, can all be automated through scripts. Therefore, we focus on two key questions: in generating docstrings and unit tests, the difference in performance between automated and manual methods.
Table 8: The average results in synthesizing compliant docstrings.
Docstring Naturalness Accuracy Clarity All Met
Programming Problem Description 97.6% 92.1% 90.9% 87.3%
Arguments and Returns - 76.4% - -
To assess the feasibility of automation, we design two experiments. First, for the canonical code of all samples in RWPB, we ask GPT-4 to generate the corresponding docstrings (i.e., the programming problem descriptions, arguments, and returns) according to the three criteria we used for annotation (i.e., naturalness, accuracy, and clarity). The prompt for generating docstrings is available in our artifact repository [54]. We then assign three experts to separately evaluate whether the docstrings generated by GPT-4 met these criteria. Table 8 presents the average ratio of compliant docstrings. From the results, we found that GPT-4 has a 97.6%, 92.1%, and 90.9% success rate in synthesizing problem descriptions that meet the “naturalness”, “accuracy”, and “clarity” criteria, respectively. The proportion of generating problem descriptions that satisfy all criteria is 87.3%. These results show that GPT-4 exhibits high accuracy in single-function comprehension and can provide an accurate problem description based on the canonical code. However, GPT-4 has room for improvement in accurately generating the functions’ arguments and returns. The success rate for accurately synthesizing arguments and returns decreases to 76.4%. For instance, GPT-4 struggles to accurately generate nested types and more complex types such as Tensors. Therefore, enhancing the LLM’s accuracy in understanding and generating function arguments and returns may become a future key research issue for automated code data synthesis.
Table 9: The average success rate in synthesizing compliant unit test inputs according to the canonical code and the docstring.
Docstring Source At Least One Execution Successful All Executions Successful
GPT-4 Generated Docstring 69.2% 66.9%
Human Written Docstring 88.6% 85.7%
Second, we evaluate GPT-4’s performance in generating inputs for unit tests given the canonical code and the docstring. We task GPT-4 to generate diverse unit test inputs, including regular and boundary cases, to cover the entire code. The docstrings are categorized into those previously generated by GPT-4 and those manually written in RWPB. The unit test output can be obtained by compiling and executing the canonical code. The experimental results are shown in Table 9. From the results, we found that when provided with human-written docstrings, the total execution success rate of GPT-4’s synthesized unit tests is 85.7%. However, when the docstrings contain inaccuracies, particularly incorrect arguments and return types, the total success rate of the synthesis decreases to 66.9%. This indicates that more accurate information must be provided to LLMs when using them to automatically generate unit tests. Additionally, we found that in the initial generation, some test samples did not achieve full coverage of the canonical code, and some raised compilation errors. To address this issue, we provide the unexecuted code segments or compilation error messages in the prompt to GPT-4 to iteratively generate additional unit tests. The results show that GPT-4 can improve the unit test coverage iteratively.
In summary, while GPT-4 ensures acceptable accuracy in automatically synthesizing programming problem descriptions, there is still room for improvement in the precision of generating the function’s arguments and returns.
7.4 Future Work
Our study used a fixed prompt template to guide LLMs in completing programming problems. This approach allows us to control experimental settings and directly investigate the types of bugs that LLMs might introduce during code generation. However, various prompts and demonstrations (i.e., in-context learning) can improve LLM’s performance on the code generation task. In the future, we plan to explore the differences in bugs introduced by LLMs when using various prompts compared to our used prompt, examining aspects such as variations in bug categories and their proportional distributions. Additionally, we ensure the uniqueness of the generated code for the same problem by setting the temperature to 0.1 and using a top-k sampling with k set to 1. In the future, we plan to investigate the effect of different decoding strategies on bug distributions.
In this study, all benchmarks we used were monolingual, meaning the canonical code was written in Python. Although recent efforts have focused on enhancing LLMs’ code generation capabilities in Python, and the results also on these benchmarks are indicative of their overall code generation proficiency, in future work, we plan to extend our methodology to multilingual scenarios, further analyzing the types and causes of bugs in code generated by LLMs across different programming languages.
We discussed a way to automate our dataset-building process, as shown in Section 7.3. We plan to develop a more effective code data synthesis pipeline that will assist not only in the evaluation process but also in the training process in the future. Finally, we introduced a self-critique-based method to enable LLMs to fix the bugs they generate. In the future, we plan to further leverage our bug categorization and experimental findings to enhance the performance of code generation and self-based bug fixing.
7.5 Threats to Validity